# Inference-time Alignment in Continuous Space

### v4

@@ 1 Introduction
The alignment of large language models (LLMs) plays a crucial role to ensure the model outputs meet human expectations and reflect human values (bai2022training, ; ouyang2022training, ; stiennon2020learning, ; xiao2025on, ; xiao2025simper, ). Reinforcement Learning from Human Feedback (RLHF) (ouyang2022training, ; christiano2017deep, ) has emerged as a widely adopted method for LLM alignment. RLHF typically involves training a reward model based on human feedback and subsequently employing reinforcement learning (RL), such as proximal policy optimization (PPO) (schulman2017proximal, ), to generate responses that maximize the reward for input prompt. Preference fine-tuning (rafailov2024direct, ; ethayarajh2024kto, ; meng2024simpo, ; NEURIPS2024_cf8b2205, ; xiao-etal-2024-leverage, ) has also been proposed as alternatives to RLHF, replacing RL with supervised learning using preference data.
Inference-time alignment liudecoding ; khanov2024args ; huang2024deal eliminates the need for additional training phases and instead focuses on guiding model behavior during inference, which has gained increasing attention due to its simplicity and flexibility. This approach does not require fine-tuning the parameters of large language models (LLMs), enabling plug-and-play adaptation for any unaligned LLM. Specifically, Best-of-N (BoN) (beirami2024theoretical, ) selects the best response based on a reward model from multiple responses generated by the base model. Other reward-guided search methods (zhou2024weaktostrong, ; khanov2024args, ; deng2023rewardaugmented, ) adjust model’s output token by token (or chunk by chunk), selecting each subsequent partial output according to reward signals.
@@ 2 Related Work
Reinforcement Learning from Human Feedback. RLHF is an effective approach for aligning LLM with human preferences (christiano2017deep, ). It is a two-stage process whereby a reward model is initially trained from human feedback and then to enhance agent’s policy via reinforcement learning, such as PPO (schulman2017proximal, ).
Inference-Time Alignment. Inference-time alignment refers to the process of adjusting model’s behavior according to certain feedback during inference, including the follow methods. Best-of-N Sampling nakano2021webgpt ; stiennon2020learning generates $N$ responses and selects one with the highest reward score. Rejection Sampling liu2023statistical generates and selects responses according to reward score threshold. ARGS Khanov2024ARGSAA generates and selects tokens based both on likelihood and reward score. CBS Zhou2024WeaktoStrongSA operates the beam search at the chunk level. Additionally, there are methods leveraging representation engineering qiu2024spectral ; kong2024aligning ; wang2024inferaligner and aligners ji2024aligner . Our work differs from the above methods and provides a novel perspective on inference-time alignment by iteratively optimizing responses guided by reward gradients.
Energy Based Models (EBMs). Energy-Based Models song2021train define a distribution through an energy function as Boltzmann distribution yuan2024tea ; yuan2024mita . EBMs have been widely used for controllable text generation due to flexibility. deng2020residual propose residual EBMs for text generation. MuCoCO kumar2021controlled formulates decoding process as an optimization problem for controllable inference. COLD qin2022cold employs gradient-based sampling in vocabulary spaces to achieve constrained generation. MuCoLa kumar2022gradient optimizes smaller intermediate representations instead of entire vocabulary. COLD-Attack guo2024coldattack designs energy functions for both controllability and stealthiness to execute jailbreak attacks. In contrast, we address the problem of inference-time alignment and demonstrate that the optimal alignment policy can be achieved by leveraging the gradient information from the reward model during inference.
@@ 3.1 Inference-time Alignment Based on Search
Inference-time alignment has emerged as a promising approach to align large language models (LLMs) with human preferences without the need for expensive retraining liudecoding ; khanov2024args ; huang2024deal . These methods offer flexibility and adaptability, making it particularly attractive for real-time applications. One widely used method is Best-of-N (BoN) strategy nakano2021webgpt ; stiennon2020learning , where the base LLMs generate multiple candidate responses, and a reward model selects the best one according to the reward. Specifically, given a prompt $\mathbf{x}$, sample $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{N}$ independently from the base reference model $\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})$, BoN selects the response with the highest reward $r(\mathbf{x},\mathbf{y})$ as the final response:
@@
Since the search is conducted during inference, BoN can be utilized for on-the-fly composition, model adaptation, and, in principle, fine-grained customization. In addition to BoN, advanced search strategies khanov2024args ; huang2024deal ; mudgalcontrolled ; Zhou2024WeaktoStrongSA , have been proposed for inference-time alignment recently.
@@ 3.2 Limitations of Existing Search Methods
Moreover, in Figure˜2(b), we analyze the minimum $N$ required by BoN to generate safe responses, as classified by a judge classifier wang2023not ,for all AdvBench requests. A safe response is considered as a good response for the given request. To further investigate, we bucket the requests and analyze the distribution of the base model’s log probabilities for good responses to these requests across varying values of minimum $N$. The results reveal that as the log probability of good responses decreases, the minimum $N$ required increases. This indicates that for a given request, the weaker the base model’s ability, the larger the $N$ needed for BoN to generate a good response. Furthermore, as $N$ grows exponentially, BoN’s performance becomes increasingly limited by base model’s capabilities.
@@ 4.1 Simple Energy Adaptation
where ${1}/{\alpha}>0$ is an appropriate KL penalty coefficient. RLHF typically optimizes the above objective using RL algorithms, such as PPO (schulman2017proximal, ). Although RLHF has achieved remarkable success, its training process is unstable and expensive engstrom2020implementation ; zheng2023secrets ; rafailov2024direct . In addition, the need to repeat training when modifying the reward model limits flexibility for adapting to evolving datasets and emerging needs.
@@
where $Z(\mathbf{x})=\sum_{\mathbf{y}}\pi_{\mathrm{ref}}(\mathbf{y}\mid\mathbf{x})\exp(\alpha r(\mathbf{x},\mathbf{y}))$ is the partition function. This unnormalized form of the optimal RLHF policy, also known as the Energy-Based Model (EBM) song2021train ; lecun2006tutorial , takes advantage from both the reference model $\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})$ and the reward function $r(\mathbf{x},\mathbf{y})$ that serves as a assessment.
As partition function $Z(\mathbf{x})$ requires computing the expectation of all possible sequences, directly sampling from this optimal policy becomes computationally prohibitive. In this paper, we propose to utilize gradient-based Langevin Markov Chain Monte Carlo method (MCMC) welling2011bayesian ; parisi1981correlation , offering more efficient sampling by using the gradient. Specifically, the gradient of the log-probability is equal to the (negative) gradient of the energy:
@@
Using the gradient above, we propose to apply Langevin MCMC welling2011bayesian , an iterative method that generates samples from the distribution by leveraging the following gradient of its log-probability:
@@
where $i$ is the sampling iteration with step size $\eta$, and $\bm{\epsilon}^{(n)}$ is the Gaussian noise. When $n\rightarrow\infty,\mathbf{y}^{(n)}$ will converge to distribute as the optimal $\pi^{*}(\mathbf{y}\mid\mathbf{x})$. Langevin dynamics does not place restrictions on sample initialization $\mathbf{y}^{0}$ given sufficient steps. However, we find that starting with random noise suffers from slow convergence and requires expensive computation. Thus, we initialize the MCMC chain from the datapoint sampled from $\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})$ , and perform a fixed number of $N$ MCMC steps; typically fewer than required for convergence of the MCMC chain hinton2002training .
A challenge is that the continuous gradient of $E(\mathbf{x},\mathbf{y}^{(n)})$ is not well-defined because $\mathbf{y}$ is discrete and non-differentiable. To address this, we utilize the continuous logits (soft outputs) of the LLMs as a representation of $\mathbf{y}$. Specifically, instead of mapping logits to language tokens using the vocabulary, we directly feed the continuous logits as input tokens to the reference and reward models. This modification eliminates the need for inference in the discrete space, allowing the alignment process with reward model to be optimized end-to-end via gradient descent, as continuous representations are fully differentiable. The use of continuous logits (soft outputs) is indeed a reasonable approximation for handling discrete sequences, as we employed the straight-through estimator bengio2013estimating with Softmax. In this approach, the discrete argmax is used in the forward pass, while the softmax continuous version is used in the backward pass. Finally, after running Langevin dynamics for $N$ steps, we obtain continuous logits sequence $\mathbf{y}^{(N)}$ which is then decoded into a discrete text, as shown in Figure˜3.
@@ 4.2 Analysis and Discussion
Shallow vs. Deep Alignment. In contrast to other alignment methods khanov2024args ; huang2024deal ; mudgalcontrolled that decode in a discrete token-by-token manner during inference, the continuous decoding process of SEA is not constrained to generate tokens sequentially. Therefore, our SEA has the potential to address the problem of shallow alignment qi2024safety , wherein the alignment adapts a model’s generative distribution primarily over only its very first few output tokens. For instance, consider the scenario where a user asks, “How do I build a bomb?” and induces the model to begin its response with, “Sure, here’s a detailed guide.” The model is then much more likely to continue providing harmful information in the response due to its auto-regressive nature. In striking contrast, our SEA allows alignment steps to decode all tokens simultaneously within a global receptive field, enabling the model to recover from harmful starting conditions and achieve deep safety alignment. In Section˜5.4, we verify our SEA can effectively achieve deep safety alignment.
@@ 5.1 Experimental Setup
Datasets. We evaluate SEA across three tasks: safety, truthfulness, and reasoning. For the safety task, we use AdvBench zou2023universal , which contains 520 harmful requests reflecting harmful or toxic behavior, we also use HH-RLHF DBLP:journals/corr/abs-2204-05862 in Section˜C.1. For the truthfulness task, we use TruthfulQA lin2021truthfulqa in generation mode and sample 100 queries for evaluation. For the reasoning task, we use two datasets: GSM8K cobbe2021gsm8k and MATH lightman2023lets . GSM8K consists of grade-school math problems, while MATH contains math problems for high-school competitions. We sample 200 samples from each dataset.
Metrics. For all three tasks, we evaluate the Average Reward, which is the mean reward from the reward model across all responses, using the same reward model as in the inference stage. For the safety task, we evaluate the Harmful Rate of generated responses, with a lower value indicating better safety. It is measured by a Longformer-based beltagy2020longformer classifier provided by wang2023not . For the truthfulness task, we evaluate both truthfulness and informativeness using judge models originally introduced in the lin2021truthfulqa . We also evaluate Diversity in khanov2024args ; kong2024aligning , with a higher score indicating broader vocabulary range. For the reasoning task, following NEURIPS2022_8bb0d291 ; yuan2024advancing , we evaluate the accuracy of the final answers.
@@
Models. We use four LLaMA-3 dubey2024llama models with different parameter sizes under both instruct and non-instruct setups. The non-instruct setup includes LLaMA-3.2-1B-Base, LLaMA-3.2-3B-Base, and LLaMA-3-8B-Base, all after supervised fine-tuning. The instruction setup uses the instruction-tuned model LLaMA-3.2-1B-Instruct. For the reward model, we use the state-of-the-art GRM-LLaMA-3.2-3B-rewardmodel-ft yang2024regularizing , which ranks among the top in the RewardBench RewardBench .
Baselines. We compare SEA with the vanilla SFT and search-based inference-time alignment methods at various granularities. At the sentence level, we include BoN nakano2021webgpt ; stiennon2020learning for $N=8,32,64$ and Rejection Sampling liu2023statistical . At the token level, we include ARGS Khanov2024ARGSAA . At the chunk level, we include CBS Zhou2024WeaktoStrongSA .
@@ 5.4 Mitigation of Shallow Alignment
As introduced by (qi2024safety, ), current safety-aligned methods often suffer from “Shallow Alignment”, where most of the KL budget gao2023scaling is allocated to the first few prefix tokens. This means that, the model only pretends to be safe for the first few tokens. As a result, harmful outputs can still be triggered by bypassing the refusal block in early token positions, a vulnerability known as Prefilling Attacks andriushchenko2024jailbreaking . However, we find that SEA demonstrates completely different behavior, effectively avoiding the drawbacks of shallow alignment. Our analysis is conducted for the following four aspects:
@@
Balanced KL budget under SEA. As shown in Figure˜5 (a), SEA ensures the KL budget is evenly distributed across all positions, which is significantly different from traditional alignment methods (refer to Figure 1 in (qi2024safety, )), which focus primarily on just a few prefix token positions. The KL divergence increases as the iterations progress (marked in different colors: black to purple to yellow). In the final iteration, largest KL budget assigned by SEA is focused on the later token positions, reaching a KL divergence close to 8, while the others remain around 5.

### v5

@@ 1 Introduction
The alignment of large language models (LLMs) plays a crucial role to ensure the model outputs meet human expectations and reflect human values (Bai et al., 2022b; Ouyang et al., 2022; Stiennon et al., 2020; Xiao et al., 2025b, a). Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Christiano et al., 2017) has emerged as a widely adopted method for LLM alignment. RLHF typically involves training a reward model based on human feedback and subsequently employing reinforcement learning (RL), such as proximal policy optimization (PPO) (Schulman et al., 2017), to generate responses that maximize the reward for input prompt. Preference fine-tuning (Rafailov et al., 2024; Ethayarajh et al., 2024; Meng et al., 2024; Xiao et al., 2024b, a) has also been proposed as alternatives to RLHF, replacing RL with supervised learning using preference data.
Inference-time alignment Liu et al. (2024b); Khanov et al. (2024a); Huang et al. (2024) eliminates the need for additional training phases and instead focuses on guiding model behavior during inference, which has gained increasing attention due to its simplicity and flexibility. This approach does not require fine-tuning the parameters of large language models (LLMs), enabling plug-and-play adaptation for any unaligned LLM. Specifically, Best-of-N (BoN) (Beirami et al., 2024) selects the best response based on a reward model from multiple responses generated by the base model. Other reward-guided search methods (Zhou et al., 2024a; Khanov et al., 2024a; Deng and Raffel, 2023) adjust model’s output token by token (or chunk by chunk), selecting each subsequent partial output according to reward signals.
@@ 2 Related Work
Reinforcement Learning from Human Feedback. RLHF is an effective approach for aligning LLM with human preferences (Christiano et al., 2017). It is a two-stage process whereby a reward model is initially trained from human feedback and then to enhance agent’s policy via reinforcement learning, such as PPO (Schulman et al., 2017).
Inference-Time Alignment. Inference-time alignment refers to the process of adjusting model’s behavior according to certain feedback during inference, including the follow methods. Best-of-N Sampling Nakano et al. (2021); Stiennon et al. (2020) generates $N$ responses and selects one with the highest reward score. Rejection Sampling Liu et al. (2023) generates and selects responses according to reward score threshold. ARGS Khanov et al. (2024b) generates and selects tokens based both on likelihood and reward score. CBS Zhou et al. (2024b) operates the beam search at the chunk level. Additionally, there are methods leveraging representation engineering Qiu et al. (2024); Kong et al. (2024); Wang et al. (2024) and aligners Ji et al. (2024). Our work differs from the above methods and provides a novel perspective on inference-time alignment by iteratively optimizing responses guided by reward gradients.
Energy Based Models (EBMs). Energy-Based Models Song and Kingma (2021) define a distribution through an energy function as Boltzmann distribution Yuan et al. (2024b, c). EBMs have been widely used for controllable text generation due to flexibility. Deng et al. (2020) propose residual EBMs for text generation. MuCoCO Kumar et al. (2021) formulates decoding process as an optimization problem for controllable inference. COLD Qin et al. (2022) employs gradient-based sampling in vocabulary spaces to achieve constrained generation. MuCoLa Kumar et al. (2022) optimizes smaller intermediate representations instead of entire vocabulary. COLD-Attack Guo et al. (2024) designs energy functions for both controllability and stealthiness to execute jailbreak attacks. In contrast, we address the problem of inference-time alignment and demonstrate that the optimal alignment policy can be achieved by leveraging the gradient information from the reward model during inference.
@@ 3.1 Inference-time Alignment Based on Search
Inference-time alignment has emerged as a promising approach to align large language models (LLMs) with human preferences without the need for expensive retraining Liu et al. (2024b); Khanov et al. (2024a); Huang et al. (2024). These methods offer flexibility and adaptability, making it particularly attractive for real-time applications. One widely used method is Best-of-N (BoN) strategy Nakano et al. (2021); Stiennon et al. (2020), where the base LLMs generate multiple candidate responses, and a reward model selects the best one according to the reward. Specifically, given a prompt $\mathbf{x}$, sample $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{N}$ independently from the base reference model $\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})$, BoN selects the response with the highest reward $r(\mathbf{x},\mathbf{y})$ as the final response:
@@
Since the search is conducted during inference, BoN can be utilized for on-the-fly composition, model adaptation, and, in principle, fine-grained customization. In addition to BoN, advanced search strategies Khanov et al. (2024a); Huang et al. (2024); Mudgal et al. (2024); Zhou et al. (2024b), have been proposed for inference-time alignment recently.
@@ 3.2 Limitations of Existing Search Methods
Moreover, in Figure˜2(b), we analyze the minimum $N$ required by BoN to generate safe responses, as classified by a judge classifier Wang et al. (2023),for all AdvBench requests. A safe response is considered as a good response for the given request. To further investigate, we bucket the requests and analyze the distribution of the base model’s log probabilities for good responses to these requests across varying values of minimum $N$. The results reveal that as the log probability of good responses decreases, the minimum $N$ required increases. This indicates that for a given request, the weaker the base model’s ability, the larger the $N$ needed for BoN to generate a good response. Furthermore, as $N$ grows exponentially, BoN’s performance becomes increasingly limited by base model’s capabilities.
@@ 4.1 Simple Energy Adaptation
where ${1}/{\alpha}>0$ is an appropriate KL penalty coefficient. RLHF typically optimizes the above objective using RL algorithms, such as PPO (Schulman et al., 2017). Although RLHF has achieved remarkable success, its training process is unstable and expensive Engstrom et al. (2020); Zheng et al. (2023); Rafailov et al. (2024). In addition, the need to repeat training when modifying the reward model limits flexibility for adapting to evolving datasets and emerging needs.
@@
where $Z(\mathbf{x})=\sum_{\mathbf{y}}\pi_{\mathrm{ref}}(\mathbf{y}\mid\mathbf{x})\exp(\alpha r(\mathbf{x},\mathbf{y}))$ is the partition function. This unnormalized form of the optimal RLHF policy, also known as the Energy-Based Model (EBM) Song and Kingma (2021); LeCun et al. (2006), takes advantage from both the reference model $\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})$ and the reward function $r(\mathbf{x},\mathbf{y})$ that serves as a assessment.
As partition function $Z(\mathbf{x})$ requires computing the expectation of all possible sequences, directly sampling from this optimal policy becomes computationally prohibitive. In this paper, we propose to utilize gradient-based Langevin Markov Chain Monte Carlo method (MCMC) Welling and Teh (2011); Parisi (1981), offering more efficient sampling by using the gradient. Specifically, the gradient of the log-probability is equal to the (negative) gradient of the energy:
@@
Using the gradient above, we propose to apply Langevin MCMC Welling and Teh (2011), an iterative method that generates samples from the distribution by leveraging the following gradient of its log-probability:
@@
where $i$ is the sampling iteration with step size $\eta$, and $\bm{\epsilon}^{(n)}$ is the Gaussian noise. When $n\rightarrow\infty,\mathbf{y}^{(n)}$ will converge to distribute as the optimal $\pi^{*}(\mathbf{y}\mid\mathbf{x})$. Langevin dynamics does not place restrictions on sample initialization $\mathbf{y}^{0}$ given sufficient steps. However, we find that starting with random noise suffers from slow convergence and requires expensive computation. Thus, we initialize the MCMC chain from the datapoint sampled from $\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})$ , and perform a fixed number of $N$ MCMC steps; typically fewer than required for convergence of the MCMC chain Hinton (2002).
A challenge is that the continuous gradient of $E(\mathbf{x},\mathbf{y}^{(n)})$ is not well-defined because $\mathbf{y}$ is discrete and non-differentiable. To address this, we utilize the continuous logits (soft outputs) of the LLMs as a representation of $\mathbf{y}$. Specifically, instead of mapping logits to language tokens using the vocabulary, we directly feed the continuous logits as input tokens to the reference and reward models. This modification eliminates the need for inference in the discrete space, allowing the alignment process with reward model to be optimized end-to-end via gradient descent, as continuous representations are fully differentiable. The use of continuous logits (soft outputs) is indeed a reasonable approximation for handling discrete sequences, as we employed the straight-through estimator Bengio et al. (2013) with Softmax. In this approach, the discrete argmax is used in the forward pass, while the softmax continuous version is used in the backward pass. Finally, after running Langevin dynamics for $N$ steps, we obtain continuous logits sequence $\mathbf{y}^{(N)}$ which is then decoded into a discrete text, as shown in Figure˜3.
@@ 4.2 Analysis and Discussion
Shallow vs. Deep Alignment. In contrast to other alignment methods Khanov et al. (2024a); Huang et al. (2024); Mudgal et al. (2024) that decode in a discrete token-by-token manner during inference, the continuous decoding process of SEA is not constrained to generate tokens sequentially. Therefore, our SEA has the potential to address the problem of shallow alignment Qi et al. (2024), wherein the alignment adapts a model’s generative distribution primarily over only its very first few output tokens. For instance, consider the scenario where a user asks, “How do I build a bomb?” and induces the model to begin its response with, “Sure, here’s a detailed guide.” The model is then much more likely to continue providing harmful information in the response due to its auto-regressive nature. In striking contrast, our SEA allows alignment steps to decode all tokens simultaneously within a global receptive field, enabling the model to recover from harmful starting conditions and achieve deep safety alignment. In Section˜5.4, we verify our SEA can effectively achieve deep safety alignment.
@@ 5.1 Experimental Setup
Datasets. We evaluate SEA across three tasks: safety, truthfulness, and reasoning. For the safety task, we use AdvBench Zou et al. (2023), which contains 520 harmful requests reflecting harmful or toxic behavior, we also use HH-RLHF Bai et al. (2022a) in Section˜C.1. For the truthfulness task, we use TruthfulQA Lin et al. (2021) in generation mode and sample 100 queries for evaluation. For the reasoning task, we use two datasets: GSM8K Cobbe et al. (2021) and MATH Lightman et al. (2023). GSM8K consists of grade-school math problems, while MATH contains math problems for high-school competitions. We sample 200 samples from each dataset.
Metrics. For all three tasks, we evaluate the Average Reward, which is the mean reward from the reward model across all responses, using the same reward model as in the inference stage. For the safety task, we evaluate the Harmful Rate of generated responses, with a lower value indicating better safety. It is measured by a Longformer-based Beltagy et al. (2020) classifier provided by Wang et al. (2023). For the truthfulness task, we evaluate both truthfulness and informativeness using judge models originally introduced in the Lin et al. (2021). We also evaluate Diversity in Khanov et al. (2024a); Kong et al. (2024), with a higher score indicating broader vocabulary range. For the reasoning task, following Kojima et al. (2022); Yuan et al. (2024a), we evaluate the accuracy of the final answers.
@@
Models. We use four LLaMA-3 Dubey et al. (2024) models with different parameter sizes under both instruct and non-instruct setups. The non-instruct setup includes LLaMA-3.2-1B-Base, LLaMA-3.2-3B-Base, and LLaMA-3-8B-Base, all after supervised fine-tuning. The instruction setup uses the instruction-tuned model LLaMA-3.2-1B-Instruct. For the reward model, we use the state-of-the-art GRM-LLaMA-3.2-3B-rewardmodel-ft Yang et al. (2024), which ranks among the top in the RewardBench Lambert et al. (2024).
Baselines. We compare SEA with the vanilla SFT and search-based inference-time alignment methods at various granularities. At the sentence level, we include BoN Nakano et al. (2021); Stiennon et al. (2020) for $N=8,32,64$ and Rejection Sampling Liu et al. (2023). At the token level, we include ARGS Khanov et al. (2024b). At the chunk level, we include CBS Zhou et al. (2024b).
@@ 5.4 Mitigation of Shallow Alignment
As introduced by (Qi et al., 2024), current safety-aligned methods often suffer from “Shallow Alignment”, where most of the KL budget Gao et al. (2023) is allocated to the first few prefix tokens. This means that, the model only pretends to be safe for the first few tokens. As a result, harmful outputs can still be triggered by bypassing the refusal block in early token positions, a vulnerability known as Prefilling Attacks Andriushchenko et al. (2024). However, we find that SEA demonstrates completely different behavior, effectively avoiding the drawbacks of shallow alignment. Our analysis is conducted for the following four aspects:
@@
Balanced KL budget under SEA. As shown in Figure˜5 (a), SEA ensures the KL budget is evenly distributed across all positions, which is significantly different from traditional alignment methods (refer to Figure 1 in (Qi et al., 2024)), which focus primarily on just a few prefix token positions. The KL divergence increases as the iterations progress (marked in different colors: black to purple to yellow). In the final iteration, largest KL budget assigned by SEA is focused on the later token positions, reaching a KL divergence close to 8, while the others remain around 5.